Destructive interference of waves occur when and where a ______ from source #1 meets a ______ from source #2.

Group of answer choices

crest, crest

crest, wavelength

trough, crest

Two identical cars travel at equal speeds but stop suddenly. The stopping force is smaller for the car with smaller stopping time.

Group of answer choices

True

False

PreviousNext

Rajon Rondo battles James Harden for a basketball. Rondo pulls leftward with 510 N of pull force, and Harden pulls rightward with 585 N of pull force. The net force on the basketball is

Group of answer choices

1095 N, rightward.

75 N, rightward.

75 N, leftward.

1095 N, leftward.
Answer #1

1. Destructive interference of waves occur when and where a trough from source #1 meets a crest from source #2.

2. Newtons second law stats Net Force on a body is the product of its mass and acceleration.

\\ \sum F = Ma \\ \text{stoping force therefore signifies deaccelration or -a} \\ \text{from kinematic equations} \\ v=u-at \\ \text{as v=0 therefore} \\ t=\frac{u}{a} \ \text{or} \ t \propto \frac{1}{a} \\ \text{Thus for a smaller stoping time t, the deacceleration a should be large} \\ \text{Therefore Larger the stopping force lesser is the stoping time.} \\ \\ \text{Therefore given statement is }\textbf{False}

3.

Force is vector, i.e it has a magnitude and a direction. In the given case since both force act in opposite direction, hence the net force is given as

\\ \vec{F}_{net}=\vec{F}_{right}-\vec{F}_{left} \\ \vec{F}_{net}=585-510\Rightarrow 75 \ \text{N} \ \text{acting towards the right}